OpenStudy (anonymous):

Gordon is evaluating some stocks he wants to purchase. You can see the function of the stock price below. The variable x represents days. 1. Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. 2. Graph the price function for Stock D. Explain the key features of the graph with complete sentences.

OpenStudy (anonymous):

3. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). 4. Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. • f(x + 2) • f(–x) • f(x) + 3

OpenStudy (anonymous):

@LiveLaughDie

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

there is no D there

OpenStudy (anonymous):

ik its asking you to create it @satellite73

OpenStudy (anonymous):

can you plz help @LiveLaughDie

OpenStudy (anonymous):

would 1 be f(x)=23(0.05)^x ? would that be okay?

OpenStudy (anonymous):

i mean f(x)=23(1.67)^x?

OpenStudy (anonymous):

what does f(x) represent?

OpenStudy (anonymous):

idk the price i think

OpenStudy (anonymous):

do you know about functions?? say you were explaining it to a friend and this friend had no idea about functions what will you tell him

OpenStudy (anonymous):

um i barely know it either i dont want to give you the wrong answer because it is very difficult for me

OpenStudy (anonymous):

that is what i would say

OpenStudy (anonymous):

well the way that i usually explain it is that f(x) in an equation lets say for example f(x)=2x+5 fx=3 lets say so you would substitute the f(x) to three and 2(3) + 5. That is kinda the only way i am able to explain it.

OpenStudy (anonymous):

you plug in x = 3 and you get out f(x) = 11 so if we said f(x) + 3, what do we mean?

OpenStudy (anonymous):

11+3?

OpenStudy (anonymous):

You would take the price and add 3 to it so effectively all of the prices bump up by $3

OpenStudy (anonymous):

o

OpenStudy (anonymous):

Sorry i gtg

OpenStudy (anonymous):

ugh okay thx

OpenStudy (anonymous):

@ProfessionalDreamer

OpenStudy (anonymous):

so sorry but i cant do this

OpenStudy (anonymous):

okay thank you